### 21 Feb 2024 v1

# What Linguistic Features and Languages are Important in LLM Translation?

Abstract
Large Language Models (LLMs) demonstrate strong capability across multiple tasks, including machine translation. Our study focuses on evaluating Llama2’s machine translation capabilities and exploring how translation depends on languages in its training data. Our experiments show that the 7B Llama2 model yields above 10 BLEU score for all languages it has seen, but not always for languages it has not seen. Most gains for those unseen languages are observed the most with the model scale compared to using chat versions or adding shot count. Furthermore, our linguistic distance analysis reveals that syntactic similarity is not always the primary linguistic factor in determining translation quality. Interestingly, we discovered that under specific circumstances, some languages, despite having significantly less training data than English, exhibit strong correlations comparable to English. Our discoveries here give new perspectives for the current landscape of LLMs, raising the possibility that LLMs centered around languages other than English may offer a more effective foundation for a multilingual model.
Keywords: Llama2, machine translation, linguistic distances
What Linguistic Features and Languages are Important in LLM Translation?
| Ryandito Diandaru, Lucky Susanto, Zilu Tang, |
| Ayu Purwarianti, Derry Wijaya |
| Bandung Institute of Technology, University of Indonesia, Boston University, |
| Monash University Indonesia |
| 13519157@std.stei.itb.ac.id, lucky.susanto@ui.ac.id, zilutang@bu.edu, ayu@itb.ac.id, |
| derry.wijaya@monash.edu |
Abstract content
1. Introduction
Large Language Models (LLMs) and their utilizations have been a popular research topic in Natural Language Processing (NLP) due to their remarkable performance on various tasks Brown et al. (2020); OpenAI (2023); Touvron et al. (2023a, b) and machine translation is no exception. Extensive evaluations Robinson et al. (2023); Hendy et al. (2023) on machine translation of the prominent GPT OpenAI (2023) models have shown that they exhibit a promising capability for high-resource machine translation although sometimes not for underrepresented languages Robinson et al. (2023); Hendy et al. (2023); Stap and Araabi (2023); Kadaoui et al. (2023).
A popular approach for low-resource NLP is to leverage other higher-resourced languages. These approaches include using them as pivot Xia et al. (2019), in transfer learning Gu et al. (2018); Nguyen and Chiang (2017) and for joint training Neubig and Hu (2018); Johnson et al. (2017). Performance gains from such methods indicate a strong influence of their presence in the training data. Given that including related languages alongside the low-resource language can often improve performance Xia et al. (2019); Poncelas and Effendi (2022); Gu et al. (2018); Nguyen and Chiang (2017); Neubig and Hu (2018); Johnson et al. (2017), it makes sense to evaluate the degree of proximity between these languages, which can be done using the vectors from the URIEL typological database Littell et al. (2017). The utilization of the URIEL database has facilitated a more guided approach to exploring the incorporation of multiple languages into a single model by leveraging linguistically-aware feature vectors from which linguistic distances can be computed. These feature vectors have been utilized by previous works in various ways including determining which language to use as transfer or pivot language Lin et al. (2019); Nambi et al. (2023) and measuring language diversity Ruder et al. (2021).
It is ideal for an LLM to be as transparent as possible to conduct a comprehensive performance evaluation. This includes details such as the specific languages, the quantity of data for each language, and the nature of the data used for training. A significant constraint in studies that assess the GPT model series Hendy et al. (2023); Robinson et al. (2023) is the fact that these models are proprietary, closed-source systems. This presents a fundamental challenge as it remains unclear which languages are included in the models. On the other hand, determining the significance of language features poses a non-trivial challenge due to its many dimensions. For instance, the URIEL database vectors encompass several linguistic features, which include FEATURAL, GENETIC, GEOGRAPHICAL, PHONOLOGICAL, INVENTORY, and SYNTACTIC distances.
In this work we are evaluating Llama2 Touvron et al. (2023b) in machine translation to highlight its multilingual capability with an emphasis on languages the model has and has not seen during training. We analyze the results against linguistic feature distances to gather insight about which features and languages in the training data are important in LLM machine translation. In summary, our contributions are as follows:
1.
We provide comprehensive evaluation on 41 languages (15 of which not seen during training) for Llama2 model families.
2.
We reveal that increasing model parameters is potentially more effective in improving translation over instruction tuning and few-shot learning.
3.
We demonstrate that syntactic similarity is not always the most important linguistic feature nor are English features always the strongest correlating factors with machine translation scores.
2. Methodology
2.1. Language Similarity Metric Evaluation
| Language | ISO 639-3 | Writing System | Genus |
Afrikaans |
afr |
Latin |
Germanic |
Galician |
glg |
Latin |
Romance |
Macedonian |
mkd |
Cyrillic |
Slavic |
Slovak |
slk |
Latin |
Slavic |
Armenian |
hye |
Armenian |
Armenian |
Basque |
eus |
Latin |
Basque |
Georgian |
kat |
Georgian |
Kartvelian |
Icelandic |
isl |
Latin |
Germanic |
Igbo |
ibo |
Latin |
Igboid |
Javanese |
jav |
Latin |
Javanese |
Sinhala |
sin |
Sinhala |
Indic |
Tagalog |
tgl |
Latin |
Greater Central Philippine |
Tamil |
tam |
Tamil |
Dravidian |
Telugu |
tel |
Telugu |
Dravidian |
Welsh |
cym |
Latin |
Celtic |
We experiment with languages reported in the training data of Llama2 Touvron et al. (2023b), the list of which and their respective ISO 639-3 codes can be found in Table 3. We refer to this set of languages as inllama. We also pick 15 languages not reported in the training data which we will refer to as outllama, presented in Table 1. It is important to highlight that languages not explicitly mentioned in Llama2 might still be present in the training data, albeit at a minuscule proportion of less than 0.005% Touvron et al. (2023b). Languages in outllama cover various language genera and writing systems. The machine translation evaluation is conducted using the FLORES-200 Guzmán et al. (2019) benchmark as it provides support for numerous low-resource languages, we exclude into-English translation directions to mitigate the risk of potential data leakage, given that the benchmark relies on Wikipedia as its English data source. We also exclude zero-shot translation as LLMs often get the language wrong in this prompting setup as reported by Robinson et al. (2023). We use a statistical machine translation metric, the BLEU Papineni et al. (2002) score and a model-based machine translation metric when applicable, the default COMET-22 Rei et al. (2022) model.
| Model | Prompt |
| Non-chat |
[SRC]: [src-sentence] |
[TGT]: [tgt-sentence] |
[SRC]: [src-sentence] |
[TGT]: |
| Chat |
This is an English to [TGT] translation, please provide the [TGT] translation for these sentences: |
[SRC]: [src-sentence] [TGT]: [tgt-sentence] |
[SRC]: [src-sentence] [TGT]: [tgt-sentence] |
Please provide the translation for the following sentence. |
Do not provide any explanations or text apart from the translation. |
[SRC]: [src-sentence] [TGT]: [tgt-sentence] |
[TGT] |
The machine translation evaluation begins with oneshot translations for both languages in inllama and outllama using the 7B model. For every language covered in this experimental setup, distances will be calculated from every language in inllama on five dimensions, namely GENETIC, GEOGRAPHICAL, INVENTORY, PHONOLOGY, and SYNTACTIC as per the URIEL typological database Littell et al. (2017). We leave out the FEATURAL distances to give each of the dimensions more focus as FEATURAL distances are combinations of all the other feature distances111For more detailed explanation of these distances, consult https://www.cs.cmu.edu/~dmortens/projects/7_project/. We calculate the Pearson correlation for each linguistic distance feature against the translation scores. We identify languages in outllama that yields under 10 BLEU score and consider them truly unseen languages. For truly unseen languages, we experiment further with model scale, chat version, and adding the shot-count (random 5 shots). For translation with chat models, we use the prompt by Robinson et al. (2023) which follows the recommendation of Gao et al. (2023) to design prompts for translation using instruction-tuned models. The prompts used on our experiments are given in Table 2
| Language | ISO 639-3 | BLEU | COMET-22 |
German |
deu |
33.68 |
0.83 |
Swedish |
swe |
37.71 |
0.87 |
Dutch |
nld |
27.45 |
0.84 |
Norwegian |
nor |
29.54 |
0.86 |
Danish |
dan |
36.21 |
0.86 |
French |
fra |
42.4 |
0.84 |
Spanish |
spa |
28.54 |
0.84 |
Italian |
ita |
28.78 |
0.85 |
Portuguese |
por |
43.21 |
0.87 |
Catalan |
cat |
35.92 |
0.84 |
Romanian |
ron |
31.58 |
0.84 |
Russian |
rus |
28.21 |
0.85 |
Polish |
pol |
22.34 |
0.83 |
Ukrainian |
ukr |
26.03 |
0.83 |
Serbian |
srp |
23.96 |
0.81 |
Czech |
ces |
24.94 |
0.82 |
Bulgarian |
bul |
29.57 |
0.83 |
Croatian |
hrv |
21.3 |
0.81 |
Slovenian |
slv |
19.51 |
0.77 |
Chinese |
zho |
19.79 |
0.82 |
Japanese |
jpn |
17.02 |
0.84 |
Vietnamese |
vie |
28.77 |
0.82 |
Korean |
kor |
11.08 |
0.78 |
Indonesian |
ind |
31.15 |
0.86 |
Finnish |
fin |
18.08 |
0.82 |
Hungarian |
hun |
18.4 |
0.78 |
| Languages in outllama | BLEU | COMET-22 |
| Afrikaans | 23.52 | 0.74 |
| Galician | 16.62 | 0.76 |
| Macedonian | 11.90 | 0.67 |
| Slovak | 11.77 | 0.68 |
| Armenian | 1.6 | 0.31 |
| Basque | 0.91 | 0.33 |
| Georgian | 2.99 | 0.31 |
| Icelandic | 2.39 | 0.35 |
| Igbo | 0.39 | 0.41 |
| Javanese | 4.33 | 0.59 |
| Sinhala | 0.25 | 0.29 |
| Tagalog | 9.65 | 0.60 |
| Tamil | 0.73 | 0.30 |
| Telugu | 0.87 | 0.33 |
| Welsh | 1.8 | 0.35 |
3. Results and Analysis
3.1. Llama2 Translation Results
One-shot 7B Llama2 translation results are presented in Table 3 and Table 4. From Table 3, we observe that none of the languages included in inllama produce a BLEU score below 10. This suggests that we can reasonably assume that Llama2 is capable of translating into all the languages it has encountered during training. However, many languages in outllama yields a BLEU score under 10, this is expected as Llama2 is presumably unfamiliar with these languages. We move forward with languages in outllama that yields a BLEU score below 10 and experiment with other variations of Llama2. We explore the effect of scale, chat version, and adding shot count and present the results on Table 5. Due to our limited compute resources we excluded the 70B and 70B-chat versions of Llama2.
Model scale enhances translation ability, but not always for instruction tuning and not so much for adding shot count.
Our experiment results presented in Table 5 demonstrate that the 13B versions of Llama2 outperforms the smaller 7B versions for all truly unseen languages. However, larger models do not seem to yield the same number of gains for every language. In best cases, 13B models increase on average as high as 2.53 BLEU with a standard deviation of 1.64. We also observed that instruction-tuning through chat versions improves translations for some languages over models with the same scale and shot count. This is observed best on Igbo and Javanese, with gains as much as 3.16 and 2.87 respectively, and worst on Tagalog, which performs worse on chat models with a decrease as big as 2.64. Adding the shot count generally improves performance although it is less drastic as model scale and instruction-tuned setups with a mean increase of 0.47 and 0.08 for non-chat and chat Llama-13B respectively. While previous three factors appears to greatly enhance Llama2’s capacity to translate into some languages, there are languages where the prospects are quite limited. For instance, for Sinhala and Tamil, scaling the model results in gains only less than 1 BLEU score and neither does using chat models or adding shot count.
| Language | 7B 1S | 7B 5S | 7B-chat 1S | 7B-chat 5S | 13B 1S | 13B 5S | 13B-chat 1S | 13B-chat 5S |
| Armenian | 1.6 | 1.95 | 2.26 | 2.43 | 2.52 | 3.03 | 2.89 | 3.03 |
| Basque | 0.91 | 1.08 | 2.98 | 3.11 | 1.52 | 1.9 | 3.72 | 3.88 |
| Georgian | 2.99 | 3.44 | 4.41 | 4.7 | 5.57 | 6.19 | 5.97 | 5.88 |
| Icelandic | 2.39 | 3.06 | 3.9 | 3.86 | 4.72 | 5.21 | 5.24 | 5.04 |
| Igbo | 0.39 | 0.59 | 1.77 | 2.04 | 0.56 | 0.67 | 3.72 | 3.49 |
| Javanese | 4.33 | 3.71 | 4.94 | 5.06 | 3.15 | 3.76 | 5.92 | 6.63 |
| Sinhala | 0.25 | 0.38 | 0.57 | 0.52 | 0.48 | 0.63 | 0.63 | 0.62 |
| Tagalog | 9.65 | 10.98 | 10.8 | 10.97 | 16.1 | 16.91 | 13.82 | 14.27 |
| Tamil | 0.73 | 1.01 | 0.82 | 1.09 | 1.79 | 2 | 1.7 | 1.56 |
| Telugu | 0.87 | 1.04 | 1.02 | 0.86 | 2.29 | 2.45 | 1.77 | 1.68 |
| Welsh | 1.8 | å 2.37 | 4.38 | 3.93 | 5.68 | 6.8 | 6.45 | 6.6 |
3.2. Language Proximity Analysis
The translation results using Llama2 7B with the one-shot prompting setup presented on Table 3 and Table 4 are used for linguistic proximity analysis. We first retrieve pre-computed distances222http://www.cs.cmu.edu/~aanastas/files/distances.zip from the URIEL database and retrieve only the distances between the languages we are translating into and the languages reported in Llama2. Self or identity distances e.g. Igbo distance to Igbo are excluded in the Pearson correlation calculation. We present our linguistic proximity analysis as heatmaps presented in Figure 1 and 2 for correlations with BLEU and COMET-22 respectively. The numbers presented are correlations between each translation language’s distance from each language in Llama2 for each feature in the URIEL database with the machine translation (MT) scores. We create several different heatmaps according to the translation languages considered. Heatmaps created are for inllama languages only, outllama languages only, both inllama and outllama languages, only Germanic languages, only Romance languages, only Slavic languages, and for languages belonging to other genus. It is important to highlight that a negative correlation implies that the closer a language is to a particular linguistic feature, the higher the MT score is likely to be and vice versa. We observe that MT score of a language correlate with Wikipedia article count of the language333Retrieved from https://meta.wikimedia.org/wiki/List_of_Wikipedias_by_language_group on October 2023, as high as 0.64 Pearson correlation with BLEU and 0.55 Pearson correlation with COMET-22.
Syntactic distance is not always the most important feature.
When including every inllama languages, BLEU and COMET-22 scores show consistently strong correlations with syntactic features, especially with Germanic and Romance languages. This finding may not be particularly groundbreaking, as we already understand that the languages in inllama predominantly belong to these language genus. However, when considering only outllama languages, translation quality seem to depend more on genetic and phonological distances although it is often inconsistent whether it is a positive or negative correlation. Syntactic distances on the other hand, show almost no correlation to MT scores except for certain languages like Norwegian and Catalan, which consistently display a strong tendency to correlate with both MT metrics. Correlation with features other than syntactic is also observed when considering languages in other genus, in which proximity with the inventory feature of Vietnamese, Dutch, German, and French are shown to correlate with COMET-22 scores.
English is not always the most syntactically important.
It is evident that when including every inllama languages, syntactic proximity with English often ranks among the top features that affects translation quality. This is also evident when considering languages from other genus; English demonstrates the most substantial syntactic correlation, although there are other languages, such as Swedish and Vietnamese, that also display some degree of correlation. However, English is often not in first place when considering languages by genus (i.e., Germanic, Slavic, and Romance) despite having the highest amount of training data. Similar to when we observe that syntactic proximities to Norwegian and Catalan have higher correlations with MT scores than syntactic proximity to English when considering only outllama languages, this phenomenon is accentuated when calculating correlations by language genus. Among Germanic languages themselves, syntactic proximity to English surprisingly shows little to no correlations with either BLEU or COMET-22 scores. Instead, Germanic languages’ MT scores appear to correlate more with syntactic proximities to Dutch, Swedish, Catalan, and Bulgarian. This is also observed with Slavic languages where the MT scores generally correlate with syntactic proximities to Germanic and Romance languages except English. The syntactic proximity to English has the lowest correlation of -0.22 with BLEU and no correlation with COMET-22 scores of the Slavic languages.
Finally, when focusing exclusively on Romance languages, it is interesting to observe that proximities to languages situated on the right side of the heatmaps i.e. other genus, exhibit higher correlations with the MT scores of Romance languages, where they show no correlation with other language subsets.
4. Related Work
Our work is in line with previous studies that assess LLMs for machine translation, resembling the work conducted by Hendy et al. (2023) and Robinson et al. (2023). We extend those evaluations by investigating the influence of the languages included in the model. This aspect of the analysis was previously unattainable in the two aforementioned papers, mainly due to the lack of transparency of the data used for training LLMs. Our method of analysis is closely related to the one included in the work of Robinson et al. (2023) where languages of interest are represented as feature vectors to do further evaluation. Our objective is to extend that exploration by encompassing other linguistic features obtained from the URIEL typological database Littell et al. (2017). We are interested in the phenomenon observed in the work of Lin et al. (2019) which also uses distances retrieved from URIEL Typological Database Littell et al. (2017). This work shows that however important dataset statistics features are compared to linguistic features, there are still cases where using them alone results in poor performance. This phenomenon drives us to conduct a more comprehensive exploration of linguistic features.
5. Conclusion
We provide a comprehensive evaluation of machine translation task in Llama2 in its seen and unseen languages. In this work, we provide machine translation scores for 26 non-English languages covered in the Llama2 model training data using the Llama2-7B model. We also added 15 additional languages not in the Llama2 training data using the 7B, 7B-chat, 13B, and 13B-chat versions. Our results show that Llama2 is capable of translating into languages it is unfamiliar with, although this phenomenon is observed only for some languages. We also demonstrate that the model scale has the most substantial impact on translation performance, while the impact of using the chat versions of Llama2 and adding the shot count varies by language. We calculate and visualize the correlations between language feature distances and machine translation scores. Our analysis reveals that syntactic similarity is not always the predominant linguistic factor influencing machine translation scores. Furthermore, despite English having the most training data, there are often other languages, whose features consistently exhibit equally or even stronger correlations in determining machine translation scores albeit having much fewer training data. Our findings pose a unique perspective on state-of-the-art LLMs, suggesting that the prevailing focus on English-centric models may not be the most optimal approach, potentially opening the door to more effective multilingual systems that could be shaped by languages other than English.
Limitations
Our research heavily depends on the language distances obtained from the URIEL typological database, as introduced by Littell et al. (2017). The original authors noted that many languages in the database may have missing features, which means the accuracy of our findings is constrained by the methods used to estimate or compensate for these missing features. Our evaluation with the COMET-22 metric is only done for languages supported in their models. However, we are aware that the model may not be equally reliable for all languages, thus the reliability of the COMET-22 heatmap goes as far as the COMET-22 model.
In an ideal scenario, it would be advantageous to include all languages from the FLORES-200 benchmark and all available versions of Llama2 to provide more evidence of the effectiveness of model scale and the overall generalizability of our findings. Unfortunately, our research is constrained by limited computational resources, preventing us from achieving this comprehensive coverage. We exclude to-English translation directions as Llama2 is likely to be trained on English Wikipedia corpora. We also exclude prompting languages in outllama using various dictionary-based prompting techniques due to the challenging work required to collect accurate dictionary entries for low-resource languages. However, we leave this for future work.
We are also aware that the chat versions of Llama2 have been intentionally trained to prevent the generation of harmful or toxic content, and this protective design may affect the quality of translations. Moreover, the chat versions of the model generate numerous artifacts in addition to the translated sentences. We have made diligent efforts to automate the output parsing process to ensure that metrics are calculated fairly. The task of human evaluation and manual parsing of the outputs is left for future work.
6. Bibliographical References
[…]
- Dryer and Haspelmath (2013) Matthew S. Dryer and Martin Haspelmath, editors. 2013. WALS Online (v2020.3). Zenodo.
- Gao et al. (2023) Yuan Gao, Ruili Wang, and Feng Hou. 2023. How to design translation prompts for chatgpt: An empirical study.
- Gu et al. (2018) Jiatao Gu, Hany Hassan, Jacob Devlin, and Victor O.K. Li. 2018. Universal neural machine translation for extremely low resource languages. In Proceedings of the 2018 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 1 (Long Papers), pages 344–354, New Orleans, Louisiana. Association for Computational Linguistics.
- Guzmán et al. (2019) Francisco Guzmán, Peng-Jen Chen, Myle Ott, Juan Pino, Guillaume Lample, Philipp Koehn, Vishrav Chaudhary, and Marc’Aurelio Ranzato. 2019. The FLORES evaluation datasets for low-resource machine translation: Nepali–English and Sinhala–English. In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP), pages 6098–6111, Hong Kong, China. Association for Computational Linguistics.
- Hendy et al. (2023) Amr Hendy, Mohamed Abdelrehim, Amr Sharaf, Vikas Raunak, Mohamed Gabr, Hitokazu Matsushita, Young Jin Kim, Mohamed Afify, and Hany Hassan Awadalla. 2023. How good are gpt models at machine translation? a comprehensive evaluation.
- Johnson et al. (2017) Melvin Johnson, Mike Schuster, Quoc V. Le, Maxim Krikun, Yonghui Wu, Zhifeng Chen, Nikhil Thorat, Fernanda Viégas, Martin Wattenberg, Greg Corrado, Macduff Hughes, and Jeffrey Dean. 2017. Google’s multilingual neural machine translation system: Enabling zero-shot translation. Transactions of the Association for Computational Linguistics, 5:339–351.
- Kadaoui et al. (2023) Karima Kadaoui, Samar M. Magdy, Abdul Waheed, Md Tawkat Islam Khondaker, Ahmed Oumar El-Shangiti, El Moatez Billah Nagoudi, and Muhammad Abdul-Mageed. 2023. Tarjamat: Evaluation of bard and chatgpt on machine translation of ten arabic varieties.
- Lin et al. (2019) Yu-Hsiang Lin, Chian-Yu Chen, Jean Lee, Zirui Li, Yuyan Zhang, Mengzhou Xia, Shruti Rijhwani, Junxian He, Zhisong Zhang, Xuezhe Ma, Antonios Anastasopoulos, Patrick Littell, and Graham Neubig. 2019. Choosing transfer languages for cross-lingual learning. In Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics, pages 3125–3135, Florence, Italy. Association for Computational Linguistics.
- Littell et al. (2017) Patrick Littell, David R. Mortensen, Ke Lin, Katherine Kairis, Carlisle Turner, and Lori Levin. 2017. URIEL and lang2vec: Representing languages as typological, geographical, and phylogenetic vectors. In Proceedings of the 15th Conference of the European Chapter of the Association for Computational Linguistics: Volume 2, Short Papers, pages 8–14, Valencia, Spain. Association for Computational Linguistics.
- Nambi et al. (2023) Akshay Nambi, Vaibhav Balloli, Mercy Ranjit, Tanuja Ganu, Kabir Ahuja, Sunayana Sitaram, and Kalika Bali. 2023. Breaking language barriers with a leap: Learning strategies for polyglot llms.
- Neubig and Hu (2018) Graham Neubig and Junjie Hu. 2018. Rapid adaptation of neural machine translation to new languages. In Proceedings of the 2018 Conference on Empirical Methods in Natural Language Processing, pages 875–880, Brussels, Belgium. Association for Computational Linguistics.
- Nguyen and Chiang (2017) Toan Q. Nguyen and David Chiang. 2017. Transfer learning across low-resource, related languages for neural machine translation. In Proceedings of the Eighth International Joint Conference on Natural Language Processing (Volume 2: Short Papers), pages 296–301, Taipei, Taiwan. Asian Federation of Natural Language Processing.
- OpenAI (2023) OpenAI. 2023. Gpt-4 technical report.
- Papineni et al. (2002) Kishore Papineni, Salim Roukos, Todd Ward, and Wei-Jing Zhu. 2002. Bleu: a method for automatic evaluation of machine translation. In Proceedings of the 40th Annual Meeting of the Association for Computational Linguistics, pages 311–318, Philadelphia, Pennsylvania, USA. Association for Computational Linguistics.
- Poncelas and Effendi (2022) Alberto Poncelas and Johanes Effendi. 2022. Benefiting from language similarity in the multilingual MT training: Case study of Indonesian and Malaysian. In Proceedings of the Fifth Workshop on Technologies for Machine Translation of Low-Resource Languages (LoResMT 2022), pages 84–92, Gyeongju, Republic of Korea. Association for Computational Linguistics.
- Rei et al. (2022) Ricardo Rei, José G. C. de Souza, Duarte Alves, Chrysoula Zerva, Ana C Farinha, Taisiya Glushkova, Alon Lavie, Luisa Coheur, and André F. T. Martins. 2022. COMET-22: Unbabel-IST 2022 submission for the metrics shared task. In Proceedings of the Seventh Conference on Machine Translation (WMT), pages 578–585, Abu Dhabi, United Arab Emirates (Hybrid). Association for Computational Linguistics.
- Robinson et al. (2023) Nathaniel R. Robinson, Perez Ogayo, David R. Mortensen, and Graham Neubig. 2023. Chatgpt mt: Competitive for high- (but not low-) resource languages.
- Ruder et al. (2021) Sebastian Ruder, Noah Constant, Jan Botha, Aditya Siddhant, Orhan Firat, Jinlan Fu, Pengfei Liu, Junjie Hu, Dan Garrette, Graham Neubig, and Melvin Johnson. 2021. XTREME-R: Towards more challenging and nuanced multilingual evaluation. In Proceedings of the 2021 Conference on Empirical Methods in Natural Language Processing, pages 10215–10245, Online and Punta Cana, Dominican Republic. Association for Computational Linguistics.
- Stap and Araabi (2023) David Stap and Ali Araabi. 2023. ChatGPT is not a good indigenous translator. In Proceedings of the Workshop on Natural Language Processing for Indigenous Languages of the Americas (AmericasNLP), pages 163–167, Toronto, Canada. Association for Computational Linguistics.
- Touvron et al. (2023a) Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro, Faisal Azhar, Aurelien Rodriguez, Armand Joulin, Edouard Grave, and Guillaume Lample. 2023a. Llama: Open and efficient foundation language models.
- Touvron et al. (2023b) Hugo Touvron, Louis Martin, Kevin Stone, Peter Albert, Amjad Almahairi, Yasmine Babaei, Nikolay Bashlykov, Soumya Batra, Prajjwal Bhargava, Shruti Bhosale, Dan Bikel, Lukas Blecher, Cristian Canton Ferrer, Moya Chen, Guillem Cucurull, David Esiobu, Jude Fernandes, Jeremy Fu, Wenyin Fu, Brian Fuller, Cynthia Gao, Vedanuj Goswami, Naman Goyal, Anthony Hartshorn, Saghar Hosseini, Rui Hou, Hakan Inan, Marcin Kardas, Viktor Kerkez, Madian Khabsa, Isabel Kloumann, Artem Korenev, Punit Singh Koura, Marie-Anne Lachaux, Thibaut Lavril, Jenya Lee, Diana Liskovich, Yinghai Lu, Yuning Mao, Xavier Martinet, Todor Mihaylov, Pushkar Mishra, Igor Molybog, Yixin Nie, Andrew Poulton, Jeremy Reizenstein, Rashi Rungta, Kalyan Saladi, Alan Schelten, Ruan Silva, Eric Michael Smith, Ranjan Subramanian, Xiaoqing Ellen Tan, Binh Tang, Ross Taylor, Adina Williams, Jian Xiang Kuan, Puxin Xu, Zheng Yan, Iliyan Zarov, Yuchen Zhang, Angela Fan, Melanie Kambadur, Sharan Narang, Aurelien Rodriguez, Robert Stojnic, Sergey Edunov, and Thomas Scialom. 2023b. Llama 2: Open foundation and fine-tuned chat models.
- Xia et al. (2019) Mengzhou Xia, Xiang Kong, Antonios Anastasopoulos, and Graham Neubig. 2019. Generalized data augmentation for low-resource translation. In Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics, pages 5786–5796, Florence, Italy. Association for Computational Linguistics.